Q.7. Calculate the potential difference between two plates when they are separated by a distance of a 0.005m and are able to hold an electron motionless between them. (Mass of electron = 9.1x10-³1


1
Expert's answer
2021-05-20T18:22:42-0400

"mg=eE"

"E=U\/d"

where "m" is mass of electron,

"e" is charge of electron,

"E" is electric field,

"U" is potential difference between two plates,

"d" is distance between two plates.


Then:

"U=\frac{mgd}{e}=\frac{9.1\cdot10^{-31}\cdot9.8\cdot0.005}{1.6\cdot^{-19}}=2.787\cdot10^{-13}\ V"